# UniSkill: Imitating Human Videos via Cross-Embodiment Skill Representations

### v3

@@ 4.1 Experimental Setup
We also compare against XSkill [6], which learns a shared skill representation to enable cross-embodiment imitation through a self-supervised learning approach. Unlike UniSkill, it requires a scene-aligned dataset, where human demonstrations are performed in the same environment and for the same task as the robot. To support this, we collect an additional $100$ human demonstrations per task to train XSkill. Note that without this additional scene-aligned human video data, XSkill fails on all tabletop tasks (i.e., 00 success). More details on the baselines are provided in Appendix C.
@@ 4.3 Cross-Embodiment Skill Representations
During pre-training on large-scale video datasets, ISD compresses a motion pattern between two frames, allowing $\mathbf{z}_{t}$ to represent a low-level skill rather than a full task. Although both UniSkill and GCBC are trained solely on demonstrations of individual tasks, we can assemble them at inference time to perform novel task combinations by leveraging the compositional nature of skills. Figure 3 (b) presents the results of task compositions in the tabletop benchmark. While GCBC fails in all evaluations, UniSkill shows robust performance across all task combinations, even with human prompts. This highlights the combinatorial nature of the learned skill representation, enabling exceptional expandability to novel tasks.
@@
To further validate scene-level generalization, we also conduct experiments in simulation, as shown in Figure 5. Even when the human prompts come from entirely different environments, UniSkill is able to successfully infer and execute the intended task. Further detail about unseen environments are provided in Appendix B.
@@
Table 1: Ablation studies evaluating the impact of datasets conducted on the LIBERO benchmark using robot prompts. Robot: Bridge and LIBERO. Human: Something-SomethingV2 and H2O.
Does UniSkill benefit from using human videos?
@@ 6 Limitations
Finally, utilizing the cross-embodiment skill representation requires fine-tuning the skill-conditioned policy to the target environment, necessitating additional action-labeled robot data. Incorporating foundational policy models may help mitigate this requirement and promote better generalization.
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2024-00333634), and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT): the Artificial Intelligence Graduate School Program, Yonsei University (RS-2020-II201361) and the Global AI Frontier Lab International Collaborative Research (No. RS-2024-00469482 & RS-2024-00509279).
@@ A.4.3 Effect of Dataset Scale and Alignment on XSkill
A.5 Ablation Studies
Effect of Dataset for Pre-training.
Effect of Depth Prediction.
Table 6(b) demonstrates the importance of incorporating depth prediction. Since our core objective is to encode dynamic information for cross-embodiment, the representation should not overly rely on semantic information, as this could lead to embodiment-specific features. As demonstrated in Figure 12(b), removing depth prediction results in a substantial drop in K-means clustering accuracy from $82.0$ to $31.7$, indicating reduced skill separation across embodiments. To mitigate this dependency, we incorporate depth prediction into ISD. When depth prediction is not used, overall performance decreases, highlighting its importance in ensuring an embodiment-agnostic skill representation.
@@ A.5 Ablation Studies
We evaluate the effectiveness of augmenting ISD inputs during policy learning. As shown in Table 6(b), eliminating this augmentation leads to large drop in performance. Because UniSkill encodes the dynamics of the video prompt, it is sensitive to changes in viewpoint or object arrangements. These results indicate that our augmentation strategy effectively mitigates these challenges.
@@
In Figure 11, we evaluate the robustness of UniSkill by varying the video speed of the human prompts. The best performance occurs at speeds of 1.00×1.$0$ and 1.25×1.25$\times$1.25 ×. Notably, performance decreases as the speed slows down. When the speed is too low, the encoded skill interval becomes short to capture meaningful action sequences.
A.6 Additional Analyses
A.6.1 Analysis of Skill-Conditioned Policy
@@ A.6.2 Analysis of Cross-Embodiment Skill
Figure 13: Visualization of the test environment and success rates across different initial towel positions. The towel’s position is shifted up to ±$\pm 12\mathrm{cm}$ from the prompt location to evaluate UniSkill’s spatial sensitivity.
@@ A.6.3 Analysis of Spatial Sensitivity of UniSkill
We select the task Push the blue towel and modify the initial position of the towel in the evaluation environment relative to its position in the prompt video. While the towel’s position is already randomized during evaluation, we extend this variation to more extreme displacements to test the limits of spatial generalization. As shown in Figure 13, the towel’s center is shifted by 00 cm, $4$ cm, $8$ cm, and $12$ cm from the original prompt position. The results show that as the position deviates further from the original, the success rate declines.
@@ B.2 Implementation Detail
For pre-training, we initialize the FSD using the InstructPix2Pix model [23] and train the ISD from scratch. For the visual encoder, we adopt the ResNet-18 [28], and for depth prediction, we utilize the pre-trained DepthAnythingV2 model [22] without further training. During pre-training, skill interval $k$ is randomly selected between $1.0s$ and $2.0s$, with the specific values determined by the frame rate of the video datasets. The image resolution is set to $256\times 256$. For policy learning, we employ diffusion policy [24] as the policy network. In real-world experiments, the policy network is pre-trained on the DROID dataset [13] and fine-tuned on the collected dataset. During both training and inference, the skill interval $k$ is fixed at $20$ frames, the image resolution is $128\times 128$, and the action dimension is set to $7$. The hyperparameters are provided in Appendix C.
@@ Appendix C Implementation Details
C.1 Skill Dynamics Modeling
@@
Denoising Step $20$
Observation Horizon $2$
@@
C.1.1 Inverse Skill Dynamics Model
@@
C.1.2 Forward Skill Dynamics Model
@@
C.2 Universal Skill-Conditioned Policy
@@ C.2.2 Simulation
C.3 Goal-conditioned Behavioral Cloning
@@ C.4 XSkill
Video Clip Length $l$ $8$
@@
Observation Horizon $2$
Observation Dimension $7$
Action Dimension $7$
@@
Since the official XSkill codebase [^1] https://github.com/real-stanford/xskill does not include complete inference code or training configurations for real-world, we re-implement the method for real-world experiments, following the paper and available codebase as closely as possible. For the reported results, we train XSkill three times and report the best performance among the runs. The training configurations used for XSkill are reported in Table 9(a) and Table 9(b).

### v4

@@ 4.1 Experimental Setup
We also compare against XSkill [6], which learns a shared skill representation to enable cross-embodiment imitation through a self-supervised learning approach. Unlike UniSkill, it requires a scene-aligned dataset, where human demonstrations are performed in the same environment and for the same task as the robot. To support this, we collect an additional $100$ human demonstrations per task to train XSkill. Note that without this additional scene-aligned human video data, XSkill fails on all tabletop tasks (i.e., 0 success). More details on the baselines are provided in Appendix C.
@@ 4.3 Cross-Embodiment Skill Representations
During pre-training on large-scale video datasets, ISD compresses a motion pattern between two frames, allowing $\mathbf{z}_{t}$ to represent a low-level skill rather than a full task. Although both UniSkill and GCBC are trained solely on demonstrations of individual tasks, we can assemble them at inference time to perform novel task combinations by leveraging the compositional nature of skills. Figure 3 (b) presents the results of task compositions in the tabletop benchmark. While GCBC fails in all evaluations, UniSkill shows robust performance across all task combinations, even with human prompts. This highlights that UniSkill captures a combinatorial space of diverse skills rather than overfitting to specific tasks, suggesting its potential scalability to a wide range of novel tasks.
@@
To further validate scene-level generalization, we also conduct experiments in simulation, as shown in Figure 5. Even when the human prompts come from entirely different environments, UniSkill is able to successfully infer and execute the intended task. Further detail are provided in Appendix B.
@@
Table 1: Ablation studies on the impact of different training datasets, conducted on the LIBERO benchmark using robot prompts. Robot: Bridge and LIBERO. Human: Something-SomethingV2 and H2O.
Does UniSkill benefit from incorporating human videos?
@@ 6 Limitations
Finally, UniSkill prefers prompt videos that exhibit robotic behaviors because its approach relies purely on motion imitation. Without semantic cues such as task descriptions, having well-aligned behaviors in the prompts enhances performance. A promising direction is to integrate UniSkill with vision-language-action (VLA) frameworks, where the VLA component provides high-level reasoning and UniSkill contributes precise motion representations for execution.
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2024-00333634), and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT): the Artificial Intelligence Graduate School Program, Yonsei University (RS-2020-II201361) and the Global AI Frontier Lab International Collaborative Research (RS-2024-00469482 & RS-2024-00509279).
@@ A.4.3 Effect of Dataset Scale and Alignment on XSkill
A.5 Ablation Studies
Effect of Dataset for Pre-training.
Effect of Depth Prediction.
Table 6(b) demonstrates the importance of incorporating depth prediction. Because our objective is to encode dynamics that transfer across embodiments, the representations should avoid overreliance on semantic cues that can entangle embodiment-specific features. This issue is amplified in real-world setting, which are more heterogeneous than simulation environments. As demonstrated in Figure 12 (b), removing depth prediction results in a substantial drop in K-means clustering accuracy from $82.0$ to $31.7$, indicating reduced skill separation across embodiments. A consistent trend appears on the Tabletop benchmark, where removing depth reduces K-means clustering accuracy from $62.8$ to $40.6$. To mitigate this dependency, ISD explicitly incorporates depth predictions. When depth prediction is not used, overall performance decreases, highlighting its importance in ensuring an embodiment-agnostic skill representation.
@@ A.5 Ablation Studies
We evaluate the effectiveness of augmenting ISD inputs during policy learning. As shown in Table 6(b), eliminating this augmentation leads to large drop in performance. Because UniSkill encodes the dynamics of the video prompt, it is sensitive to changes in scene color and object appearance. These results indicate that our augmentation strategy effectively mitigates these challenges. We use colorjittering for the augmentation.
@@
In Figure 11, we evaluate the robustness of UniSkill by varying the video speed of the human prompts. The best performance occurs at speeds of $1.00\times$ and $1.25\times$. Notably, performance decreases as the speed slows down. When the speed is too low, the encoded skill interval becomes short to capture meaningful action sequences.
@@
A.6 Additional Analyses
A.6.1 Analysis of Skill-Conditioned Policy
@@ A.6.2 Analysis of Cross-Embodiment Skill
Figure 13: Visualization of the test environment and success rates across different initial towel positions. The towel’s position is shifted up to $\pm 12\mathrm{cm}$ from the prompt location to evaluate UniSkill’s spatial sensitivity.
@@ A.6.3 Analysis of Spatial Sensitivity of UniSkill
We select the task Push the blue towel and modify the initial position of the towel in the evaluation environment relative to its position in the prompt video. While the towel’s position is already randomized during evaluation, we extend this variation to more extreme displacements to test the limits of spatial generalization. As shown in Figure 13, the towel’s center is shifted by 0 cm, $4$ cm, $8$ cm, and $12$ cm from the original prompt position. The results show that as the position deviates further from the original, the success rate declines.
@@ B.2 Implementation Detail
For pre-training, we initialize the FSD using the InstructPix2Pix model [23] and train the ISD from scratch. For the visual encoder, we adopt the ResNet-18 [28], and for depth prediction, we utilize the pre-trained DepthAnythingV2 model [22] without further training. During pre-training, skill interval $k$ is randomly selected between $1.0s$ and $2.0s$, with the specific values determined by the frame rate of the video datasets. The image resolution is set to $256\times 256$. For policy learning, we employ diffusion policy [24] as the policy network. In real-world experiments, both UniSkill and GCBC are pre-trained on the DROID dataset [13] and fine-tuned on the collected dataset. During both training and inference, the skill interval $k$ is fixed at $20$ frames, the image resolution is $128\times 128$, and the action dimension is set to $7$. The hyperparameters are provided in Appendix C.
@@ Appendix C Implementation Details
C.1 Inference Procedure
UniSkill’s policy is trained to produce actions conditioned on skill representations. At inference, given a video prompt, it derives a sequence of skill representations from the video using a sliding-window procedure. Specifically, let the frames be $I_{1},\dots,I_{p}$ and fixed interval $k$. For each $t=1,\dots,p-1$, we form the pair $(I_{t},\;I_{\min(t+k,\,p)})$ and extract a skill $z_{t}$, so we obtain $z_{1},\dots,z_{p-1}$. For the last k steps we use $I_{p}$ as the future frame. For each timestep $t$, $z_{t}$ and observation $obs_{t}$ is then fed into the policy, which outputs the action executed in the environment.
C.2 Skill Dynamics Modeling
@@
Denoising Step 2020
Observation Horizon 22
@@
C.2.1 Inverse Skill Dynamics Model
@@
C.2.2 Forward Skill Dynamics Model
@@
C.3 Universal Skill-Conditioned Policy
@@ C.3.2 Simulation
C.4 Goal-conditioned Behavioral Cloning
@@ C.5 XSkill
Video Clip Length $l$ 88
@@
Observation Horizon 22
Observation Dimension 77
Action Dimension 77
@@
Since the official XSkill codebase [^1] https://github.com/real-stanford/xskill does not include complete inference code or training configurations for real-world, we re-implement the method for real-world experiments, following the paper and available codebase as closely as possible. Unless otherwise noted, we keep the hyperparameters from the original XSkill setup without modifications, including the number of skill prototypes. For the reported results, we train XSkill three times and report the best performance among the runs. The training configurations used for XSkill are reported in Table 9(a) and Table 9(b).